# LlamaRestTest: Effective REST API Testing with Small Language Models

### v1

@@ 3.3.2. Parameter-Efficient Fine-Tuning
Fine-tuning plays a pivotal role in enhancing the performance of LlamaREST-IPD and LlamaREST-EX by allowing the models to adapt to specific tasks, such as inter-parameter dependency (IPD) detection and example value extraction. Fine-tuning can be approached in two ways: Full Parameter Fine-Tuning (FFT) and Parameter-Efficient Fine-Tuning (PEFT). While FFT requires large datasets and extensive computational resources, PEFT is designed to work efficiently with smaller datasets, making it a better fit for our task.
Since our IPD dataset contains only 551 parameters, it is insufficient for FFT, as prior research has shown that fewer than 1,000 examples do not yield significant performance improvements with FFT (Vieira et al., 2024). To address this, we employed PEFT, which optimizes the model by fine-tuning only a subset of its parameters. This selective updating process allows it to efficiently adapt to new tasks using smaller datasets and fewer computational resources.
The two most commonly used PEFT methods are LoRA (Low-Rank Adaptation)(Hu et al., 2021) and QLoRA(Dettmers et al., 2024). Both approaches allow fine-tuning with significantly reduced computational requirements. LoRA focuses on injecting trainable low-rank matrices into transformer layers, while QLoRA quantizes model weights during training, further reducing memory usage and computational costs. Given the limitations of our compute environment, which includes a Google Colab machine with a 40GB A100 GPU, we opted for QLoRA. QLoRA’s quantization compresses the model, allowing it to efficiently fit within our memory constraints while still performing the necessary updates.
We based our fine-tuning configuration on the recommended settings from Meta’s official guidelines (Meta, 2024a), making necessary adjustments to accommodate our hardware. Specifically, we reduced the batch size to 8 to fit within the 40GB VRAM of the A100 GPU. The fine-tuning process leverages a structured dataset specifically designed for inter-parameter dependency (IPD) detection and example value extraction tasks, using custom tokens to represent the relationships between parameters. These tokens guide the model’s understanding of the dependencies and example values, allowing it to update only relevant parts of the model during fine-tuning. This targeted adaptation allows for faster convergence and lower computational costs compared to full parameter tuning.
@@ 3.3.3. Quantization
After fine-tuning, quantization was applied to optimize the model for deployment in resource-constrained environments and improve speed. Quantization reduces the model’s memory footprint by representing its weights with fewer bits, making it faster and more efficient while maintaining acceptable performance. Using the llama.cpp library, we experimented with three popular quantization levels to balance model size and accuracy: 2-bit, 4-bit, and 8-bit quantization (Frantar et al., 2022; Yao et al., 2021; Shen et al., 2020). Each level presents different trade-offs between memory usage and performance, with 2-bit offering the highest compression and 8-bit maintaining nearly full accuracy.
These quantized versions allowed us to assess the efficiency-effectiveness trade-offs for various deployment scenarios in REST API testing. The 2-bit quantization is ideal for highly memory-constrained devices like edge environments (e.g., 2.96 GB), 4-bit strikes a balance for mid-tier resources (e.g., 4.58 GB), and 8-bit is suited for relatively high-resource environments where performance is a priority (e.g., 7.95 GB). By offering these different quantized versions of the model, we aimed to assess the trade-offs between efficiency (model size and speed) and effectiveness (accuracy) for various deployment scenarios in REST API testing.
@@ 3.4. REST API Testing
Figure 2 illustrates the LlamaRestTest framework, which performs adaptive REST API testing by integrating the ARAT-RL (Adaptive REST API Testing with Reinforcement Learning) approach (Kim et al., 2023b) with fine-tuned models, LlamaREST-IPD and LlamaREST-EX. The process begins with the Specification Parser, which extracts key elements from the OpenAPI specification, including available operations and required parameters. Based on this, ARAT-RL initializes a Q-Table that prioritizes operations and parameters by assigning higher values to more frequently used resources, indicating their relative importance.
@@
When repeated failures occur (e.g., multiple 4xx responses), LlamaREST-IPD and LlamaREST-EX are activated. The LlamaREST-IPD model uses the same prompt employed during fine-tuning:
@@
This prompt directs the model to analyze parameter descriptions and server response messages, identifying dependencies, such as when a parameter is required only if another is present.
Similarly, LlamaREST-EX generates example values for parameters using the following prompt:
@@
Once parameters and values are refined, the Request Sender submits the API request. Depending on the response, the Q-Table Updator adjusts the priorities of the operation-parameter combinations. Successful requests lower the priority of that combination, encouraging exploration, while failures assign positive rewards, driving further investigation into problematic areas.
@@ 4. Evaluation
    RQ1: How do LlamaREST-EX and LlamaREST-IPD compare with the state-of-the-art REST API testing assistant tool, RESTGPT, in terms of identifying inter-parameter dependencies and generating semantically valid testing inputs for REST API testing?
@@
    RQ5: How do parameter description, server response messages, LlamaREST-IPD, and LlamaREST-EX contribute to the overall performance of LlamaRestTest in terms of code coverage and operation coverage?
@@
REST Service Lines of Code #Operations
@@
Market 7945 13
Project-Tracking-System 3784 67
YouTube-Mock 2422 1
@@
Ohsome API 122
Table 2. Internal Server Errors Collected Over 10 Runs
@@
Market 10 10 10 10 10
Project-Tracking-System 10 10 10 10 10
YouTube-Mock 0 0 0 0 0
@@ 4.1. Experiment Setup
Our experimental framework was set up on an M1 MacBook Pro running Sonoma 14.4.1 with 64GB of memory. We installed all the services from our benchmark suite, along with the REST API testing tools we intended to evaluate. To prevent any dependency conflicts, we restored the database for each service before running each test. Each service and tool was hosted with default settings and tested sequentially to avoid interference. We continuously monitored CPU and memory usage to ensure that resource limitations did not impact the performance of the testing tools. Each tool was tested for one hour, a standard duration in the field (Kim et al., 2022; Zhang and Arcuri, 2023; Kim et al., 2023b; Liu et al., 2022; Golmohammadi et al., 2023), to provide a fair comparison. To minimize the effect of randomness on the results, we repeated each test 10 times and averaged the outcomes except for the ablation study (RQ5 used the 3 runs setting).
@@
For a thorough evaluation, we used the same REST API testing tools and services employed by ARAT-RL (Kim et al., 2023b), the most recent REST API testing tool available. ARAT-RL also features the largest benchmark dataset among the tools we reviewed. Specifically, MoRest (Liu et al., 2022) evaluated six services, RESTler tested one service, and EvoMaster evaluated five services, while ARAT-RL’s benchmark dataset included 10 RESTful services. These tools are used in ARAT-RL work because of the popularity and performance, so we used the same set of tools, but with the latest versions. [^1] We used the latest release version or the latest commit when a release version was unavailable: RESTler v9.2.4, EvoMaster v3.0.0, ARAT-RL v0.1, and MoRest—obtained directly from the authors, as the website was unavailable.
@@
Additionally, we incorporated services from the RESTGPT study due to the availability of a ground truth for extracted rules (Kim et al., 2023c). Out of the 16 services, we excluded SCS and NCS, as they were authored by EvoMaster’s developers, to avoid potential bias. We also excluded OCVN due to authentication issues and omitted OMDB, a toy online service with only one API operation that all tools could process in seconds. Ultimately, we used 12 services: Features-Service, Language-Tool, REST-Countries, Genome-Nexus, Person-Controller, User-Management, Market, Project-Tracking-System, OhSome, YouTube-Mock, and Spotify. Table 2 provides details on these open-source services, including their lines of code and the number of API operations.
@@
The key metrics for evaluating effectiveness and error detection included code coverage, the number of successfully processed operations in the specification, and the number of 500 status codes, which are among the most common metrics (Golmohammadi et al., 2023). To gather code coverage data, we used Jacoco (Team, 2023). The number of processed operations and the number of errors (e.g., 500 status codes) were both collected using scripts from the ARAT-RL repository (Kim”, 2024). This script tracks 500 status codes for each operation and removes duplicated 500 errors with the server response message.
@@ 4.2. RQ1: Comparison with RESTGPT
Table 3. Comparison of semantically valid parameter value generation among LlamaREST-EX, Llama3-8B, and RESTGPT. ’FT’ indicates fine-tuning, and ’Q’ indicates quantization.
@@
Language-Tool 22.58% 58.82% 64.52% 69.23% 82.98% 60.00%
Spotify 73.74% 73.81% 75.86% 81.71% 66.67% 0.00%
YouTube-Mock 14.75% 67.27% 70.31% 73.08% 73.33% 21.67%
Genome-Nexus 10.66% 35.29% 35.48% 57.02% 36.25% 9.00%
@@
Table 4. Comparison of inter-parameter dependency rule generation performance among LlamaREST-IPD, RESTGPT, and Llama3-8B. ’FT’ indicates fine-tuning, and ’Q’ indicates quantization. ’O’ indicates correct rule generation, and ’-’ indicates no correct generation.
@@
YouTube-Mock eventType IF eventType THEN type=video - O O O O O
@@
To assess the effectiveness of LlamaREST-IPD in identifying inter-parameter dependency (IPD) rules and LlamaREST-EX in generating semantically valid testing inputs, we compared both models with RESTGPT, a state-of-the-art tool for enhancing specifications. The comparison utilized the RESTGPT dataset, which includes parameter names, descriptions, inter-parameter dependency rules, and the corresponding values generated by RESTGPT. Additionally, we evaluated Llama3-8B, the base model for both LlamaREST-IPD and LlamaREST-EX, along with various quantization variants (2-bit, 4-bit, and 8-bit).
@@
Table 3 presents seven services from the RESTGPT benchmark, alongside the precision of semantically valid value generation for LlamaREST-EX, RESTGPT, and Llama3-8B. A recent study proposed evaluating value generation based on semantic validity, a method also adopted by RESTGPT: a semantically valid input must be coherent with the API domain (Alonso et al., 2023). For instance, ”Berlin” is a valid input for the parameter addressLocality in DHL’s API, whereas ”dog” is not. For RESTGPT, we used the generated values from its artifact, while for Llama3-8B, we applied the same prompts used by RESTGPT. Precision was determined by two external developers independently assessing whether the generated values were semantically valid, with any discrepancies resolved through discussion until a consensus was reached.
@@
Llama3-8B, while competitive on some APIs, generally underperforms with an average accuracy of just 24.94%. It struggles significantly on APIs like Spotify (0%) and Ohsome (2.78%), underscoring its limitations in generating semantically valid values. However, fine-tuning the model, as demonstrated with LlamaREST-EX, leads to significant performance improvements.
@@
LlamaREST-EX, especially in its FT version, shows substantial gains over Llama3-8B, achieving an average accuracy of 72.44%. For example, on Ohsome, the FT version of LlamaREST-EX reaches 92.16%, outperforming RESTGPT’s 76.83%, and on Genome-Nexus, it scores 57.02%, surpassing RESTGPT’s 36.25%. Overall, LlamaREST-EX with fine-tuning outperforms RESTGPT, which achieved 68.82%, compared to LlamaREST-EX’s 72.44%. Even the quantized versions of LlamaREST-EX (4-bit and 8-bit) show reasonable performance, with 60.21% and 66.12% precisions, respectively.
@@
Table 4 presents the available inter-parameter dependency rules identified through parameter descriptions for three services among the seven services in Table 3. Llama3-8B demonstrates limited capability in detecting IPD rules, identifying only 2 rules across all services. LlamaREST-IPD, particularly in its fine-tuned (FT) version, significantly outperforms both Llama3-8B and RESTGPT. The FT variant of LlamaREST-IPD identifies 12 correct IPD rules, compared to RESTGPT’s 9, highlighting its ability to capture more complex relationships. Even the quantized versions of LlamaREST-IPD (4-bit and 8-bit) demonstrate consistent improvements in performance, while the 2-bit version identified only 1 correct IPD rule.
@@
Fine-tuning and quantization of LlamaREST models significantly improve the accuracy of semantically valid input generation and inter-parameter dependency rule detection, outperforming both the base model and state-of-the-art RESTGPT.
4.3. RQ2: Impact of Fine-Tuning and Quantization in REST API Testing

### v2

@@ 3.3.2. Parameter-Efficient Fine-Tuning
Fine-tuning plays a pivotal role in enhancing the performance of LlamaREST-IPD and LlamaREST-EX by enabling the models to adapt to the tasks of IPD detection and example value generation. Fine-tuning can be performed in two ways: Full Parameter Fine-Tuning (FFT) and Parameter-Efficient Fine-Tuning (PEFT). FFT requires large datasets and extensive computational resources, whereas PEFT is designed to work efficiently with smaller datasets, making it a better fit for our application. Our IPD dataset contains only 551 parameters, which is insufficient for FFT, as prior research has shown that fewer than 1,000 examples do not yield significant performance improvements with FFT (Vieira et al., 2024). We therefore employed PEFT, which optimizes the model by fine-tuning only a subset of its parameters. This selective updating process allows it to efficiently adapt to new tasks using smaller datasets and fewer computational resources.
The two most commonly used PEFT methods are LoRA (Hu et al., 2022) and QLoRA (Dettmers et al., 2023). Both approaches allow fine-tuning with significantly reduced computational requirements. LoRA focuses on injecting trainable low-rank matrices into transformer layers, whereas QLoRA quantizes model weights during training, further reducing memory usage and computational costs. Given the limitations of our compute environment, which includes a Google Colab machine with a 40GB A100 GPU, we opted for QLoRA. QLoRA’s quantization compresses the model, letting it fit within our memory constraints while still performing the necessary updates.
We based our fine-tuning configuration on the recommended settings from Meta’s official guidelines (Meta, 2024a), making necessary adjustments to accommodate our hardware. Specifically, we reduced the batch size to 8 to fit within the 40GB VRAM of the A100 GPU. As mentioned, the fine-tuning process leverages a structured dataset specifically designed for IPD detection and value generation tasks, using custom tokens to represent the relationships between parameters. These tokens guide the model’s understanding of the dependencies and example values, allowing for updates of only the relevant parts of the model during fine-tuning. This adaptation allows for faster convergence and lower computational costs compared to full parameter tuning.
@@ 3.3.3. Quantization
After fine-tuning, we applied quantization to optimize the model for deployment in resource-constrained environments and improve speed. Quantization reduces the model’s memory footprint by representing its weights with fewer bits, making it faster and more efficient while maintaining acceptable performance. Using the llama.cpp library, we experimented with three popular quantization levels to balance model size and accuracy: 2-bit, 4-bit, and 8-bit quantization (Frantar et al., 2023; Yao et al., 2021; Shen et al., 2020). Each level presents different trade-offs between memory usage and performance, with 2-bit offering the highest compression and 8-bit maintaining nearly full accuracy.
These quantized versions let us assess the trade-off between effectiveness (accuracy) and efficiency (model size and inference speed) for various deployment scenarios in REST API testing. The 2-bit quantization is ideal for highly memory-constrained devices such as edge environments (e.g., 2.96 GB), 4-bit strikes a balance for mid-tier resources (e.g., 4.58 GB), and 8-bit is suited for relatively high-resource environments where performance is a priority (e.g., 7.95 GB).
@@ 3.4. REST API Testing
Figure 2 illustrates the LlamaRestTest framework, which integrates the ARAT-RL approach for adaptive REST API testing. A key component of ARAT-RL is the Q-Table, which prioritizes API operations and parameters. The Q-Table Updater updates Q-values based on responses received from API requests. Specifically, successful responses (2xx) reduce the Q-value, encouraging exploration of other operations, while failure responses (4xx or 500) increase the Q-value, directing further testing toward problematic areas. This update follows the Q-learning formula:
(1) $Q(s,a)\leftarrow Q(s,a)+\alpha[r+\gamma\max Q(s^{\prime},a^{\prime})-Q(s,a)]$
where $\alpha$ is the learning rate, $\gamma$ is the discount factor, and $r$ is the reward assigned based on the response. As shown in Figure 2, the Request Sender interacts with the Q-Table Updater, dynamically adjusting priorities to optimize testing coverage.
@@
When repeated failures occur (e.g., multiple 4xx responses), LlamaREST-IPD and LlamaREST-EX are activated. The LlamaREST-IPD model is prompted in the same format employed during fine-tuning:
@@
This prompt directs the model to analyze parameter descriptions and server response messages, identifying dependencies, such as when a parameter is required only if another parameter is present.
Similarly, LlamaREST-EX generates example values for parameters using the following prompt structure, which was used for fine-tuning:
@@
After parameters and values are refined, the Request Sender submits the API request. Depending on the API response, the Q-Table Updater adjusts the priorities of the operation-parameter combinations. Successful requests lower the priority of that combination, encouraging exploration, whereas failures assign positive rewards, driving further probing of problematic parts of the API.
@@ 4. Evaluation
    RQ1: How do LlamaREST-EX and LlamaREST-IPD compare with the state-of-the-art REST API testing assistant tool, RESTGPT, in terms of identifying inter-parameter dependencies and generating valid inputs for REST API testing?
@@
    RQ5: How do analysis of parameter description, analysis of server response messages, LlamaREST-IPD, and LlamaREST-EX contribute to the overall performance of LlamaRestTest in terms of code coverage and operation coverage?
@@
REST Service Lines of Code Operations
@@
Market 7945 13
Person-Controller 601 12
Project-Tracking-System 3784 67
@@
YouTube 2422 1
@@
Table 2. Internal server errors detected over 10 runs.
@@
Market 10 10 10 10 10
Person-Controller 80 80 80 80 80
Project-Tracking-System 10 10 10 10 10
@@
YouTube 0 0 0 0 0
@@ 4.1. Experiment Setup
Our experiments were conducted on an M1 MacBook Pro running Sonoma 14.4.1 with 64GB of memory. We installed all the services from our benchmark suite, along with the REST API testing tools we intended to evaluate. To prevent any dependency conflicts, we restored the database for each service before running each test. Each service and tool was hosted with default settings and tested sequentially to avoid interference. We continuously monitored CPU and memory usage to ensure that resource limitations did not impact the performance of the testing tools. Each tool was run for one hour, a standard time budget used in the area (Kim et al., 2022; Zhang and Arcuri, 2023; Kim et al., 2024a; Liu et al., 2022; Golmohammadi et al., 2023), to provide a fair comparison. To minimize the effect of randomness on the results, we repeated each test 10 times and averaged the outcomes. [^2] For the ablation study (RQ5), we performed only three runs due to its extensive evaluation time.
@@
For our evaluation, we relied on the same set of REST API testing tools and services used by ARAT-RL (Kim et al., 2024a). Accordingly, we compared AutoRestTest with ARAT-RL (Kim et al., 2024a), EvoMaster (Arcuri, 2019), MoRest (Liu et al., 2022), and RESTler (Atlidakis et al., 2019). Specifically, we used the latest released version or the latest commit when a release was unavailable: RESTler v9.2.4, EvoMaster v3.0.0, ARAT-RL v0.1, MoRest (obtained directly from the authors).
@@
The ARAT-RL benchmark dataset consists of 10 RESTful services. In addition to these services, we included services from the RESTGPT study (Kim et al., 2024b) due to the availability of a ground truth of extracted rules. Out of the total 16 services, we excluded SCS and NCS because they were written by EvoMaster’s authors, and we aimed to avoid potential bias. We also excluded OCVN due to authentication issues. Lastly, we excluded OMDB, which is a toy online service with only one API operation that all testing tools can process in a second. Ultimately, we used 12 services: Features Service, Language Tool, REST Countries, Genome Nexus, Person Controller, User Management Microservice, Market Service, Project Tracking System, Ohsome, [^3] Ohsome has an open-source version (GIScience, 2025a), but we used the online version (GIScience, 2025b) as the open-source version consistently produced server errors. YouTube-Mock, and Spotify. Table 2 provides details on these services, including the lines of code (for the open-source services) and the number of API operations.
@@
The key metrics for evaluating the effectiveness and fault-detection ability included code coverage, the number of successfully processed operations in the specification, and the number of 500 status codes encountered, which are among the most commonly used metrics (Golmohammadi et al., 2023). To measure code coverage, we used JaCoCo (Team, 2023). The number of processed operations and the number of errors detected (500 status codes) were collected using a script from the ARAT-RL repository (Kim, 2024). That script tracks 500 status codes for each operation and removes duplicate 500 errors based on server response messages.
@@ 4.2. RQ1: Comparison with RESTGPT
Table 3. Comparison of semantically valid parameter value generation by LlamaREST-EX, Llama3-8B, and RESTGPT (‘FT’ indicates fine-tuning and ‘Q’ indicates quantization).
@@
YouTube 14.75% 67.27% 70.31% 73.08% 73.33% 21.67%
@@
Table 4. Comparison of inter-parameter dependency rule generation performance among LlamaREST-IPD, RESTGPT, and Llama3-8B (‘FT’ indicates fine-tuning, ‘Q’ indicates quantization, ‘O’ indicates correct rule generation, and ‘-’ indicates failure to generate the rule).
@@
YouTube eventType IF eventType THEN type=video - O O O O O
@@
To assess the effectiveness of LlamaREST-IPD in identifying IPD rules and LlamaREST-EX in generating semantically valid testing inputs, we compared both models with RESTGPT, a state-of-the-art tool for enhancing specifications. We also evaluated Llama3-8B, the base model for both LlamaREST-IPD and LlamaREST-EX, along with various quantization variants (2-bit, 4-bit, and 8-bit). To measure effectiveness, we used as ground truth, the RESTGPT dataset of extracted rules, including IPDs and example values (Kim et al., 2024b). The RESTGPT dataset contains rules manually extracted from the OpenAPI specifications of nine services, but IPD and value constraint rules occur for only three and seven of the services, respectively; for the other services, these rules were not described in their OpenAPI specifications. Thus, our evaluation of value generation is based on seven of the services from the RESTGPT benchmark, whereas the evaluation of IPD rule extraction is based on three of the services.
Table 3 presents the results on the precision of semantically valid value generation for LlamaREST-EX, RESTGPT, and Llama3-8B for seven services. Table 4 further narrows the focus to three of these services, presenting the available IPD rules identified through parameter descriptions.
@@
We evaluated value generation based on semantic validity, as proposed by a recent study and adopted by RESTGPT: a semantically valid input must be coherent with the API domain (Alonso et al., 2023). For instance, “Berlin” is a valid input for the parameter addressLocality in DHL’s API, whereas “dog” is not. For RESTGPT, we used the generated values from its artifact; for Llama3-8B, we applied the same prompts used by RESTGPT. Precision was determined by two external developers independently assessing whether the generated values were semantically valid, with any discrepancies resolved through discussion to reach a consensus.
@@
Llama3-8B, although competitive on some APIs, generally performs poorly, with an average accuracy of just 24.94%. It struggles significantly on APIs such as Spotify (0%) and Ohsome (2.78%), underscoring its limitations in generating semantically valid values. However, fine-tuning the model, as demonstrated with LlamaREST-EX, leads to significant performance improvement.
@@
LlamaREST-EX, especially its FT version, shows substantial gains over Llama3-8B, achieving an average accuracy of 72.44%. For example, on the Ohsome API, the FT version of LlamaREST-EX reaches 92.16%, outperforming RESTGPT’s 76.83%; similarly, on Genome-Nexus, it scores 57.02%, achieving considerably higher precision than RESTGPT (36.25%). Overall, LlamaREST-EX with fine-tuning outperforms RESTGPT, which achieved 68.82% accuracy, compared to LlamaREST-EX’s 72.44%. Even the quantized versions of LlamaREST-EX (4-bit and 8-bit) show reasonable performance, with 60.21% and 66.12% precision, respectively.
LlamaREST-EX generally performs well, but there are two cases where it is less effective than RESTGPT: Language-Tool and FDIC. These results highlight some limitations of our approach and provide insights for future improvements. For Language-Tool, the low performance is primarily due to LlamaREST-EX’s difficulty in handling complex parameter constraints, particularly for region-specific formats. For example, with the altLanguage parameter, LlamaREST-EX output “en” instead of the correct “en-GB”. RESTGPT, leveraging its larger model size, was able to correctly interpret the format specification from the parameter description. In the case of FDIC, the lower performance came from the specialized nature of the API, which requires more domain-specific financial knowledge that our fine-tuning process did not fully capture.
These cases highlight the trade-off between model size and performance. Although our approach with LlamaRestTest demonstrates that fine-tuning a smaller, open-source model can significantly enhance task-specific effectiveness and make it competitive with much larger models in many cases, there are still scenarios where larger models can have an advantage in handling very specific or complex constraints.
@@
Table 4 presents the IPD rules identified through parameter descriptions for three services among the seven services in Table 3. Llama3-8B demonstrates limited capability in detecting IPD rules, identifying only two rules across all services. LlamaREST-IPD, particularly in its fine-tuned version, significantly outperforms both Llama3-8B and RESTGPT. The FT variant of LlamaREST-IPD identifies 12 correct IPD rules, compared to nine by RESTGPT, highlighting its ability to capture more complex relationships. Even the quantized versions of LlamaREST-IPD (4-bit and 8-bit) demonstrate consistent improvements in performance, but the 2-bit version has considerable performance degradation, identifying only one IPD rule correctly.
@@
Fine-tuning significantly improves the accuracy of semantically valid input generation and IPD rule detection, outperforming both the base Llama3-8b model and the state-of-the-art RESTGPT tool. Moreover, with 8-bit and 4-bit quantization, the fine-tuned model remains competitive with RESTGPT, while still performing considerably better than the base model.
Figure 3. Code coverage achieved on 10 runs for open-source services using LlamaRestTest configured with the base Llama3-8B model, the fine-tuned model (FT), and the quantized models (Q).
4.3. RQ2: Impact of Fine-Tuning and Quantization on REST API Testing